# openCFS: Open Source Finite Element Software for Coupled Field Simulation - Part Acoustics

### v2

@@ 1 Introduction
Its flexible meshing and interfacing between different field equations using coupling terms is the core strength of openCFS. Over the last years, the module acoustic (openCFS-Acoustics) and the coupling to the mechanic field (vibroacoustic [6]) and fluid dynamics as a source term (aeroacoustics [45, 61, 37]) was one of the foci. For instance, vibroacoustic simulations are performed using full coupling inside openCFS, solving the acoustics and solid mechanical field in parallel [67], and the coupling is realized via coupling matrices (this idea is described in [20]). Another possibility is to prescribe the surface normal velocity as a boundary condition and perform a forward coupling simulation using the acoustic module only [72]; this is only recommended in special occasions where feedback is negligible (e.g., sound radiation from a stiff, closed mechanical structure). In contrast, some cases exist where the acoustics have strong back-coupling properties onto the solid mechanical field [34, 29]. In aeroacoustics, usually, the segregated workflow is applied [42] where in a separate flow simulation using a state-of-the-art flow solver is calculating the flow solution [50] and then integrated by the module openCFS-Data [57] onto the acoustic grid as aeroacoustic source terms that are then propagated using an acoustic simulation. The particular focus of interest is numerical methods to solve these coupled fields and connected applications. The convergence of simulations using non-conforming interfaces for stationary and rotating domains [17, 23, 64, 36] have been performed in [63]. The particular strength is the strong connection to engineering research and the regular use of the software for source code and software solution development in industrial and basic research projects. It is hosted on GitLab (gitlab.com/openCFS), with the accompanying package API pyCFS, the automated build environment, the testing examples, and the documentation. The software includes several physical fields
@@ 2.1 Standard wave equation
Regarding the definition of the wave equation on the domain $\Omega$, the Dirichlet boundary $\Gamma_{\mathrm{D}}$, and the Neumann boundary $\Gamma_{\mathrm{N}}$ (see Fig. 1), we derive the weak formulation of the wave equation.
@@ 2.2.1 Frequency dependent speed of sound (time-harmonically)
with a frequency-dependent equivalent bulk modulus $K_{\mathrm{eq}}(\omega)$ and density $\rho_{\mathrm{eq}}(\omega)$. This type of wave equation is used to model various absorber materials for acoustic treatment [3, 9].
@@ 2.2.3 Spatially dependent speed of sound
$\ddot{p}_{\mathrm{a}}-\nabla\cdot\left(c^{2}(T)\nabla p_{\mathrm{a}}\right)=0\,,$ (3)
@@ 2.3 Pressure/Velocity form of the wave equation
The advection form of the wave equation in terms of the pressure and particle velocity, as unknown, is provided in the AcousticMixedPDE method. Today, no active development is carried out here.
@@ 2.3.1 Viscous acoustics (Linearized Navier-Stokes Equation)
In the method LinFlowPDE, the linearized Navier-Stokes equations are implemented. For example, one can look at [13, 33].
@@ 2.4 Postprocessing of acoustic fields (acousticSplitPDE)
A general compressible fluid field can be decomposed into a longitudinal and transversal process using the acousticSplitPDE [59, 58, 51, 61, 38].

### v3

@@ 1 Introduction
Its flexible meshing and interfacing between different field equations using coupling terms is the core strength of openCFS. Over the last years, the module acoustic (openCFS-Acoustics) and the coupling to the mechanic field (vibroacoustic [6], with examples see [^1] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Coupledfield/MechAcou) and fluid dynamics as a source term (aeroacoustics [45, 61, 37]) was one of the foci. For instance, vibroacoustic simulations are performed using full coupling inside openCFS, solving the acoustics and solid mechanical field in parallel [67], and the coupling is realized via coupling matrices (this idea is described in [20]). Another possibility is to prescribe the surface normal velocity as a boundary condition and perform a forward coupling simulation using the acoustic module only [72]; this is only recommended in special occasions where feedback is negligible (e.g., sound radiation from a stiff, closed mechanical structure). In contrast, some cases exist where the acoustics have strong back-coupling properties onto the solid mechanical field [34, 29]. In aeroacoustics, usually, the segregated workflow is applied [42] where in a separate flow simulation using a state-of-the-art flow solver is calculating the flow solution [50] and then integrated by the module openCFS-Data [57] onto the acoustic grid as aeroacoustic source terms that are then propagated using an acoustic simulation. The particular focus of interest is numerical methods to solve these coupled fields and connected applications. The convergence of simulations using non-conforming interfaces for stationary and rotating domains [17, 23, 64, 36] (e.g. example with rotating interface [^2] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Acoustics/RotationNCAcouPotential) have been performed in [63]. The particular strength is the strong connection to engineering research and the regular use of the software for source code and software solution development in industrial and basic research projects. It is hosted on GitLab (gitlab.com/openCFS), with the accompanying package API pyCFS, the automated build environment, the testing examples, and the documentation. The software includes several physical fields
@@ 2.1 Standard wave equation
Regarding the definition of the wave equation on the domain $\Omega$, the Dirichlet boundary $\Gamma_{\mathrm{D}}$, and the Neumann boundary $\Gamma_{\mathrm{N}}$ (see Fig. 1), we derive the weak formulation of the wave equation. Examples can be found here [^3] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Acoustics and for a first example we recommend [^4] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Acoustics/Abc2d.
@@ 2.2.1 Frequency dependent speed of sound (time-harmonically)
with a frequency-dependent equivalent bulk modulus $K_{\mathrm{eq}}(\omega)$ and density $\rho_{\mathrm{eq}}(\omega)$. This type of wave equation is used to model various absorber materials for acoustic treatment [3, 9]. An example simulation for an acoustic meta-material placed in a duct can be found in the openCFS-Testsuite  [^5] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Acoustics/Muffler1MPPnmg.
@@ 2.2.3 Spatially dependent speed of sound
$\ddot{p}_{\mathrm{a}}-\nabla\cdot\left(c^{2}(T)\nabla p_{\mathrm{a}}\right)=0\,.$ (3)
An example simulation with a temperature-dependent PML can be found in the openCFS-Testsuite  [^8] https://gitlab.com/openCFS/Testsuite/-/blob/master/TESTSUIT/Singlefield/Acoustics/WaveFlowTempPml.
@@ 2.2.4 Convective wave equation
An example simulation and usage can be found in the openCFS-Testsuite  [^9] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Acoustics/PCWE.
@@ 2.3 Pressure/Velocity form of the wave equation
The advection form of the wave equation in terms of the pressure and particle velocity, as unknown, is provided in the AcousticMixedPDE method. Today, no active development is carried out here. Example simulations and usage can be found in the openCFS-Testsuite  [^10] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/AcousticMixed.
@@ 2.3.1 Viscous acoustics (Linearized Navier-Stokes Equation)
In the method LinFlowPDE, the linearized Navier-Stokes equations are implemented. For example, one can look at [13, 33]. Example simulations and usage can be found in the openCFS-Testsuite  [^11] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/LinFlow.
@@ 2.4 Postprocessing of acoustic fields (acousticSplitPDE)
A general compressible fluid field can be decomposed into a longitudinal and transversal process using the acousticSplitPDE [59, 58, 51, 61, 38]. Example simulations and usage can be found in the openCFS-Testsuite  [^12] https://gitlab.com/openCFS/Testsuite/-/tree/master/TESTSUIT/Singlefield/Split.